OpenStudy (anonymous):

The Brown's are renting a home and are planning to set up their trash service. They are considering two local companies. Company A charges a $50 setup fee and $15 per month. This can be represented by the linear function C = 15x + 50. Company B charges $100 for the setup fee and $12 per month. This can be represented by the linear function C = 12x + 100. They are planning to rent their home and use trash service for 6 months. Which statement is true?

OpenStudy (texaschic101):

what are your statements ?

OpenStudy (anonymous):

The cost of using Company B is the same as the cost of using Company A. The cost of using Company B is $32 less than the cost of using Company A. Eliminate The cost of using Company A is $32 less than the cost of using Company B. The cost of using Company B is $18 less than the cost of using

OpenStudy (texaschic101):

company A: C = 15(6) + 50 = 90 + 50 = 140 company B : C = 12(6) + 100 = 72 + 100 = 172 172 - 140 = 32 ANSWER : The cost of using company A is $32 less then using company B.

OpenStudy (anonymous):

thank you! would you mind helping me with another problem please?

OpenStudy (texaschic101):

I can try

OpenStudy (anonymous):

The prices of tickets for the local rock concert were $16 for Friday night and $28 for Saturday night. The total attendance for the two nights was 1090. If the total receipts were $24,232, how many people atten saturday?

OpenStudy (texaschic101):

16f + 28s = 24,232 s + f = 1090 --> f = 1090 - s now sub 1090 - s in for f in the 1st equation 16(1090 - s) + 28s = 24,232 17440 - 16s + 28s = 24,232 17440 + 12s = 24,232 12s = 24,232 - 17440 12s = 6792 s = 6792/12 s = 566 s + f = 1090 566 + f = 1090 f = 1090 - 566 f = 524 Saturday's attendance = 566 Friday's attendance = 524

OpenStudy (anonymous):

Thank you! You are awesome!

OpenStudy (texaschic101):

glad to help :)